OpenStudy (anonymous):

help will fan and medal a.Inconsistent, no solution b.Inconsistent, infinitely many solutions c.Consistent, one solution d.Consistent , infinitely many solutions question below

OpenStudy (anonymous):

\[\left\{ 2x+y =2 \right\}\]

OpenStudy (anonymous):

\[\left\{ 6x+3y=12 \right\}\]